the base of a right prism is an equilateral triangle whose edges is 10cm each and its lateral edge is 80 cm. Determine the total surface area of the prism in cm2


1
Expert's answer
2022-05-17T06:47:12-0400

Total surface area of a prism:

"S=2B+Ph"

Where

B = Base area

P = perimeter of the base

h = height of the prism

Since we have equilateral triangle, its area is:

"B=\frac {\sqrt3} 4 a^2"

and perimeter is "P=10*3=30 cm"

where a is the edge of triangle.

So, the total surface area of a prism:

"S=\frac {2\sqrt3} 4 *10^2+30*80\approx2486,6 cm^2"